A manager wants to build 2-sigma x-bar control limits for a process. The target value for the mean of the process is 10 units, and the variance of the process is 36. If 9 units are to be taken in each sample, what will be the lower and upper control limits, respectively?

Select one:

a. [2,18]

b. [6,16]

c. [8.66,11.34]

d. [8,12]

e. [6,14]
Answer #1


1. E. 6, 14


MEAN / X-BAR = 10
STANDARD DEVIATION = 6 = SQRT(36)
Z = 2
SAMPLE SIZE(N) = 9

UCL = X-BAR + (Z * STDEV / SQRT(N)) = 10 + (2 * 6 / SQRT(9)) = 14
LCL = X-BAR - (Z * STDEV / SQRT(N)) = 10 - (2 * 6 / SQRT(9)) = 6

** Leaving a thumbs-up would really help me out. Let me know if you face any problems.